# Local Mechanisms of Compositional Generalization in Conditional Diffusion

### v2

@@ E.1 CLEVR Dataset, architecture, and training details
Our experiments cover two different conditioning setups. Location-conditioning conditions on 2D object locations, implemented as an 2D integer array representing a grid_size × grid_size grid over the image recording the count of objects whose center falls within the grid cell, as shown in Figure 9. The count is usually 0 or 1 but can be greater than 1 if object centers happen to land within the same grid cell. We take grid_size=16 in all experiments. We either record the locations of all objects, or just one of the objects (randomly selected), in the conditioning grid, depending on the experiment. Color conditioning is implemented as a 8-dimensional integer array (there are 8 possible colors) indicating the count of objects with the corresponding color. In all experiments we condition only a single attribute (either location or color) at a time, with all other attributes sampled randomly and not conditioned on.
@@
In all experiments, the model is trained with a batch size of $2048$ over $128\times 2^{20}$ samples, repeating samples if needed. Our training procedure is identical to EDM2 Karras et al. (2024) except that we do weight renormalization after the weights are updated. At inference, we use raw conditional diffusion scores, without applying any guidance/CFG (Ho & Salimans, 2022).
@@ E.2 Length generalization evaluation
In Table 1, we define $K_{\max}$ as the maximum value such that the model “succeeds” for every $1\leq K\leq K_{\max}$, with “success” defined for $1\leq K\leq N$ as generating exactly $K$ objects, and for $K>N$ as generating at least K2K-2 and $K$ objects more than 25% of the time, with objects appearing in approximately correct locations and with acceptable image quality. This metric is intentionally generous as it is intended to capture the largest $K$ for which the model “sometimes succeeds”, rather than requiring perfect performance. To assess these criteria, we manually count over 8 samples of each composition. We test on up to $K=12$ locations and up to $K=8$ colors. Table 2 provides a similar analysis for additional experiments with other conditioning configurations such as labeling a random number of objects.
@@
In Table 1 and Figure 8 we observe improvements in length-generalization as we increase the maximum number $M$ of objects the model was trained on, for models that length-generalize at all (i.e. Experiments 1 and 3). We note that for larger numbers of locations $K$ at inference-time, the objects become crowded (less independent). We hypothesize that in general, models trained on 1M1-M objects could learn to represent clusters of 1M1-M objects, as well as how to compose multiple clusters. (This is still consistent with our theory: it is a conditional projective composition where the conditioners are subsets. If the underlying data has this type of compositional structure, a trained model could learn to group individual conditioners into subsets in order to exploit it.) If this is the case, a model trained on 1-3 objects could generate, for instance, 12 objects, by composing 4 clusters of 3 objects each. This would mean that models trained on more objects (larger $M$) could more easily length-generalize to higher $K$ where objects become crowded.
Figure 10: Locality metrics for CLEVR length generalization for Experiments 1, 2, and 3, conditioned OOD on 4 locations, as in Figure 2. (Left) Conditioner gradient magnitudes at selected pixels (loc 0–3 and center), with colors indicated individual conditioners. Experiment 1 shows strong conditional-locality at high noise (each pixel responds only to its corresponding conditioner), whereas Experiment 2 exhibits weaker, non-local responses. Experiment 3 has conditional-locality explicitly enforced by the architecture. (Middle) Pixel gradient area required to cover 90% of score gradient “energy” (sum of squared magnitudes). Insets illustrate the selected square regions at loc 2 at a few timesteps. We observe similar pixel-locality between Experiment 1 and Experiment 2; the pixel gradients are highly localized at both high and low noise, but delocalize during intermediate timesteps. Experiment 3 has pixel-locality explicitly enforced by the architecture. (Right) Total pixel gradient energy (over entire image), which is higher at low noise levels, consistent with conditioners dominating the score field at high noise and pixel interactions emerging later in denoising.
E.3 Details of Experiment 3: Local Patch-Based Architecture Intervention.
@@
For Experiment 3 (Figure 1), we developed a local variant of the EDM2 model that processes images as a grid of overlapping patches. The image is divided into a grid_size × grid_size grid of cells, where each cell has size m = resolution / grid_size. We set grid_size = 16 to match the location-conditioning grid. For each grid cell at position $(i,j)$, we extract a patch of size $M=(2k+1)\times m$, where $k$ is the neighborhood radius. We used $k=2$ for the experiments in this paper. Each patch is conditioned only on the location-conditioners that fall within the patch – that is, a $(2k+1)\times(2k+1)$ subgrid of the full conditioning grid. Each patch is also conditioned on its absolute location (that is, the model is not equivariant). This is implemented by appending the patch center coordinates $i,j$ to the flattened location-conditioner grid to form the complete conditioner. The training procedure uses a patch sampling approach balancing positive and negative examples for efficiency. For each training image, we randomly sample two patches: one “positive” patch with at least one active conditioner, and one “negative” patch with no active conditioners. Standard EDM2 loss is applied to each patch and losses are averaged.
@@
In Experiment 3, we trained the local model on the same dataset and location-conditioning (labeling only a single object) as in Exp. 2, showing that Exp. 3 restores length-generalization that failed in Exp. 2. As a test, we also verified that setting $k=8$ (which for grid size 16 makes the patches the size of the image) reproduces the behavior of Exp. 2 (i.e. length-generalization fails). We also trained the local model on a dataset with only a single object per image and show that it length-generalizes up to 6 objects in this case, whereas a standard model trained on only one object per image only generates one object per image at test time, regardless of conditioning (Figure 4).
E.4 Detail of color-conditioned CLEVR feature-space study.
In Figure 5, the similarity matrices show cosine similarities (heuristic Equation 3) between network layer activations, when the model is conditioned on different colors (that is, the $i,j$ entry is the cosine similarity between the mean difference vectors for colors $i$ and $j$). We perform the study on MCScale layers, which play a similar role to cross-attention in EDM2. Figure 5 shows the encoder MCScale layers only; Figure 11 provides the complete length-generalization plots and cosine similarities between colors for all EDM2 MCScale layers. We identify a possible disentangled feature-space within the early encoder layers: specifically the layers with resolution 128 and 64 (in the figure, the average of these layers is labeled the “feature-space”). Note that this is an empirical and subjective identification, based on our observation that the cosine-similarity heuristic suggests disentanglement within these particular layers; we currently have no theoretical basis for predicting whether and where such disentanglement might occur within the network. Nevertheless, the fact that disentanglement does seem to appear anywhere within the network’s internal representation helps to explain the observed partial length generalization according to our theory.
Figure 11: Additional detail for Figure 5 color-conditioned CLEVR study. Complete length-generalization plots (1 to 8 colors; >5>5 colors is OOD) and cosine similarities between colors for all EDM2 MCScale layers (both encoder and decoder).
Figure 12: Pixel locality structure across time for CIFAR-10, using the same EDM2 architecture from the CLEVR experiments. (Right) Each column represents a timestep $t$, and the top row shows the predicted denoised images via learned scores. Lower rows show heatmaps of the pixel gradient magnitude (average absolute Jacobian from an selected pixel to all other pixels). Evaluation at three pixel locations confirms prior empirical observations of Kamb & Ganguli (2024); Niedoba et al. (2024) that the effective local neighborhood size is large at high noise levels. (Left) Locality metrics as in Figure 10; details in Appendix F.
@@ Appendix F Pixel- and Conditional-Locality Gradient Analyses
In this section we describe the pixel- and conditional-locality analyses used in Figures 2, 6, 12, 10.
@@
$G_{i,j}(x,y,t)=\frac{1}{B}\sum_{b=1}^{B}\sum_{c=1}^{3}\left|\frac{\partial\widehat{x}_{0}[b,c,y,x]}{\partial x_{t}[b,c,i,j]}\right|$ (9)
@@
$I_{k}(x,y,t)=\left\|f_{\theta}(x_{t},t,C)[x,y]-f_{\theta}(x_{t},t,C_{-k})[x,y]\right\|_{2}$ (10)
@@
In Figure 10 we directly plot the gradient influence $I_{k}(x,y)$ for each conditioner $k$ across all times $t$. We quantify spatial locality by finding the smallest square centered at each target location that contains a specified fraction (90%) of the total gradient energy, i.e. the smallest square $s$ s.t. $\sum_{(i,j)\in\text{square}_{s}}G_{i,j}^{2}\geq 0.9\times\sum_{i,j}G_{i,j}^{2}$.
@@
Figure 13: Additional detail for Figure 7. Cosine similarities between selected concepts for all SDXL cross-attention layers. The mid block activations at low and mid noise levels are also shown; note that the disentanglement is weaker at low noise. Disentanglement is most evident in the early mid blocks at high noise, suggesting these layers may serve as a compositional feature-space.
For Figure 6 we use a pretrained SDXL model (out-of-the-box, no finetuning), specifically stabilityai/stable-diffusion-xl-base-1.0, with the prompt “a beautiful photograph with a horse in the middle, a dog on the left, and a cat on the right” – which contains implicit location conditioning. We perform a locality analysis similar to the one we used for CLEVR as described in Appendix F, but with a few adaptations. The gradient computation is performed in the VAE latent space. Since SDXL uses more complex conditioning, the conditioning influence analysis takes $C=\{\text{text\_embeds},\text{pooled\_embeds},\text{time\_ids}\}$ and $\mathcal{C}_{-k}$ represents the conditioning with word $k$ removed from the text prompt. Specifically, we first first tokenize the input prompt into individual words, remove common words like “a”, “the”, etc., and for each remaining word creating a modified prompt by removing that word. We then compute the score difference w.r.t. the modified prompt. We rank the words by their influence magnitude to study how specific words affect the score at particular pixel locations. We consider a set of words “dominant” if their minimum influence magnitude is at least 2x the influence of the next-ranked word. We ran the model with 50 inference steps using the standard EulerDiscreteScheduler stopped at steps $t=1,25,50$.

### v3

@@ E.1 CLEVR Dataset, architecture, and training details
Our experiments cover a few different conditioning setups. Grid-style location-conditioning conditions on 2D object locations, implemented as an 2D integer array representing a grid_size × grid_size grid over the image recording the count of objects whose center falls within the grid cell, as shown in Figure 6. The count is usually 0 or 1 but can be greater than 1 if object centers happen to land within the same grid cell. We take grid_size=16 in all experiments. We either record the locations of all objects, or just one of the objects (randomly selected), in the conditioning grid, depending on the experiment.
List-style location conditioning also conditions on 2D object locations, but lists the (embedded) xy-locations of each object in an array padded with enough slots for up to 10 objects (with each location placed in a randomly chosen slot).
Color conditioning is implemented as a 8-dimensional integer array (there are 8 possible colors) indicating the count of objects with the corresponding color. In all experiments we condition only a single attribute (either location or color) at a time, with all other attributes sampled randomly and not conditioned on.
@@
In all experiments, the model is trained with a batch size of $2048$ over $128\times 2^{20}$ samples, repeating samples if needed. Our training procedure is identical to EDM2 Karras et al. (2024) except that we do weight renormalization after the weights are updated. At inference, we use raw conditional diffusion scores, without applying any guidance/CFG (Ho and Salimans, 2022).
Figure 12: Locality metrics for CLEVR length generalization for Experiments 1, 2, 3, and 2L, conditioned OOD on 4 locations, as in Figure 2. (Left) Conditioner gradient magnitudes at selected pixels (loc 0–3 and center), with colors indicating individual conditioners. Experiment 1 shows strong conditional-locality at high noise (each pixel responds only to its corresponding conditioner), whereas Experiments 2 and 3 exhibit non-local responses. Experiment 2L has conditional-locality explicitly enforced by the architecture. (Middle) Pixel gradient area required to cover 90% of score gradient “energy” (sum of squared magnitudes). Insets illustrate the selected square regions at loc 2 at a few timesteps. We observe similar pixel-locality between Experiments 1 and 2; the pixel gradients are highly localized at both high and low noise, but delocalize during intermediate timesteps. Experiment 3 exhibits pixel non-locality even at high noise. Experiment 2L has pixel-locality explicitly enforced by the architecture. (Right) Total pixel gradient energy (over entire image), which is higher at low noise levels, consistent with conditioners dominating the score at high noise and pixel interactions emerging later in denoising.
@@ E.2 Length generalization evaluation
In Table 1, we define $K_{\max}$ as the maximum value such that the model “sometimes succeeds” for every $1\leq K\leq K_{\max}$, with “success” defined as generating $K$ objects at least 50% of the time, and at least K2K-2 objects at least 90% of the time. This metric is intentionally generous as it is intended to capture the largest $K$ for which the model “sometimes succeeds”, rather than requiring perfect performance. To assess these criteria in Table 1, we use a ResNet-18 classifier (99.74% test accuracy) trained to count 1-10 CLEVR objects at specified locations, evaluated over 1024 samples per configuration. Note that the classifier counts the total number of objects, so for models that draw “extra” objects at non-conditioned locations (e.g. Exp. 2, which was trained with only one location labeled but draws additional objects), $K_{\max}$ may slightly overestimate the number of objects at correct conditioned locations. Nevertheless, this does not materially affect the conclusions: Exp. 2 still fails to consistently place objects at conditioned locations beyond $\sim$3, even when trained on 1-5 objects. We test on up to $K=12$ locations. In Table 2 we provide the complete counts for all experiments shown in Table 1. Table 3 provides a similar analysis for additional experiments with other conditioning configurations such as labeling a random number of objects.
@@
In Table 1 and Figure 10 we observe improvements in length-generalization as we increase the maximum number $M$ of objects the model was trained on, for models that length-generalize at all (i.e. Experiments 1 and 2L). We note that for larger numbers of locations $K$ at inference-time, the objects become crowded (less independent). We hypothesize that in general, models trained on 1M1-M objects could learn to represent clusters of 1M1-M objects, as well as how to compose multiple clusters. (This is still consistent with our theory: it is a conditional projective composition where the conditioners are subsets. If the underlying data has this type of compositional structure, a trained model could learn to group individual conditioners into subsets in order to exploit it.) If this is the case, a model trained on 1-3 objects could generate, for instance, 12 objects, by composing 4 clusters of 3 objects each. This would mean that models trained on more objects (larger $M$) could more easily length-generalize to higher $K$ where objects become crowded.
Experiment $K_{\max}$ Automated classifier counts
0 1 2 3 4 5 6 7 8 9 10
Exp.1, trained 1 1 0 1021 3 0 0 0 0 0 0 0 0
Exp.1, trained 1-3 7 0 0 0 0 0 3 205 808 8 0 0
Exp.1, trained 1-5 10 0 0 0 0 0 0 0 0 8 327 689
Exp.2, trained 1 1 0 1021 3 0 0 0 0 0 0 0 0
Exp.2, trained 1-3 2 0 342 387 293 2 0 0 0 0 0 0
Exp.2, trained 1-5 3 0 175 267 208 185 168 21 0 0 0 0
Exp.3, trained 1-3 3 0 63 170 790 1 0 0 0 0 0 0
Exp.3, trained 1-5 6 0 0 0 0 6 606 367 45 0 0 0
Exp.2L, trained 1 5 0 0 0 0 262 653 107 2 0 0 0
Exp.2L, trained 1-3 9 0 0 0 0 0 0 2 63 291 351 317
Exp.2L, trained 1-5 10 0 0 0 0 0 0 2 11 66 189 756
Exp.3L, trained 1-3 9 0 0 0 0 0 15 64 187 316 228 214
Table 2: $K_{\max}$ counts for Table 1 experiments, evaluated using the automated classifier over 1024 samples. The “Len” column indicates the number of locations conditioned on (equal to $K_{\max}$); columns represent number of objects generated at conditioned locations, and rows contain counts of images that contain the number of objects listed in the column.
Table 3: Additional location-conditioning experiments (Figure 2). $K_{\max}$ values and count distributions for additional conditioning variants beyond Table 1: All labels means every object was labeled (as in Exp. 1), Single label means only one object (randomly selected) was labeled (as in Exp. 2), and Rand num labels means a random number of objects were labeled. The “Color” column indicates the color used in Figure 2. $K_{\max}$ is evaluated using the automated classifier over 1024 samples at the final training epoch.
Experiment Color $K_{\max}$ Automated classifier counts
0 1 2 3 4 5 6 7 8 9 10
All labels, trained 1 brown 1 0 1021 3 0 0 0 0 0 0 0 0
All labels, trained 1-2 orange 5 0 0 0 69 373 580 2 0 0 0 0
All labels, trained 1-3 red 7 0 0 0 0 0 3 205 808 8 0 0
All labels, trained 1-5 blue 10 0 0 0 0 0 0 0 0 8 327 689
Single label, trained 1-3 cyan 2 0 342 387 293 2 0 0 0 0 0 0
Rand labels, trained 1-2 purple 3 0 0 13 1009 2 0 0 0 0 0 0
Rand labels, trained 1-3 green 5 0 0 0 0 151 873 0 0 0 0 0
E.2.1 Figure 2 Details
In Figure 2 plots conditional locality vs. length generalization for a variety of models. Each model is shown in a different color (with different shapes indicating different stages during training). The early, mid, and late stages are 16777216, 33554432, 134217728 images seen, respectively. The color assignments, conditioning variants, and $K_{\max}$ values for each experiment are given in Table 3.
In Figure 2 (Right), the $x$-axis shows length-generalization, defined as the number of locations to which the model can generalize beyond the number on which it was trained (e.g. +4 for a model trained on 1-3 locations that generalizes to 7). The number of locations to which the model can generalize is evaluated via $K_{\max}$ as described in Section E.2. The complete counts are given in Table 3. The conditional locality ($y$-axis) metric is described in Appendix F.
E.3 Details of Experiment 2L: Local Patch-Based Architecture Intervention.
@@ E.2.1 Figure 2 Details
For Experiment 2L (Figure 1 and 8), we developed a local variant of the EDM2 model that processes images as a grid of overlapping patches. The image is divided into a grid_size × grid_size grid of cells, where each cell has size m = resolution / grid_size. We set grid_size = 16 to match the location-conditioning grid. For each grid cell at position $(i,j)$, we extract a patch of size $M=(2k+1)\times m$, where $k$ is the neighborhood radius. We used $k=2$ for the experiments in this paper. Each patch is conditioned only on the location-conditioners that fall within the patch – that is, a $(2k+1)\times(2k+1)$ subgrid of the full conditioning grid. Each patch is also conditioned on its absolute location (that is, the model is not equivariant). This is implemented by appending the patch center coordinates $i,j$ to the flattened location-conditioner grid to form the complete conditioner. The training procedure uses a patch sampling approach balancing positive and negative examples for efficiency. For each training image, we randomly sample two patches: one “positive” patch with at least one active conditioner, and one “negative” patch with no active conditioners. Standard EDM2 loss is applied to each patch and losses are averaged.
@@
In Experiment 2L, we trained the local model on the same dataset and location-conditioning (labeling only a single object) as in Exp. 2, showing that Exp. 2L length-generalizes while Exp. 2 fails. As a test, we also verified that setting $k=8$ (which for grid size 16 makes the patches the size of the image) reproduces the behavior of Exp. 2 (i.e. length-generalization fails). We also trained the local model on a dataset with only a single object per image and show that it length-generalizes up to 5 objects in this case, whereas a standard model trained on only one object per image only generates one object per image at test time, regardless of conditioning (Figure 8).
E.4 Experiment 3L
Figure 13: (Left) Experiment 3L applies a causal intervention to the failing Exp. 3: we modify the model architecture to explicitly enforce local conditional scores, use the same training data (1-3 objects) and conditioning as Exp. 3, and find that Exp. 3L length-generalizes while Exp. 3 failed. Note that both Exp. 3 and Exp. 3L struggle somewhat with location-accuracy even in-distribution, and this inaccuracy persists out-of-distribution for Exp. 3L. Nevertheless, Exp. 3L displays much stronger length-generalization than Exp. 3, enabling much higher object counts. (Right) Length generalization in color-conditioned CLEVR. We test length generalization in a CLEVR model conditioned on colors (rather than locations), trained on 1-5 objects. Each row is conditioned on the first $K$ colors in order (blue, brown, cyan, gray, green, purple, red, yellow), e.g. row 4 is conditioned on blue, brown, cyan, gray. We test up to 8 colors (>5>5 colors is OOD), and find generalization up to 7 colors, suggesting that compositional structure may exist in a learned feature-space.
Analogous to Experiment 2L, we apply a local causal intervention to the non-length-generalizing Experiment 3 and show that it enables length-generalization, as shown in Figure 13. In Experiment 3L, the model denoises individual patches conditioned only on the location-conditioners that fall within the patch, with conditioners represented as a list as in Experiment 3. The setup is as described in Section E.3, with the only difference being the list-style conditioner. Specifically, the original conditioner lists the locations of each object in an array padded with enough slots for up to 10 objects (with each object placed in a randomly-chosen slot); the patch conditioner includes only the conditioned locations that fall within the current patch, re-centered relative to the patch center, and placed within the padded array in their original random slots. Each patch is also conditioned on its absolute location as described in Section E.3. We find that this model struggles with location accuracy even in-distribution (up to 3 objects) and this inaccuracy persists out-of-distribution object counts. Nevertheless the local model of Exp. 3L still displays much stronger length-generalization than the original model of Exp. 3: when trained on 1-3 objects, the original model struggles to ever produce more than 3 objects at any locations, while the local model can draw at least 12 objects (though not always at the desired locations). For Exp. 3 trained on 1-3 objects, we have $K_{\max}=3$, while for Exp. 3L $K_{\max}=9$ (note that $K_{\max}$ requires approximately correct locations so it is lower than the maximum number of objects the model can actually produce in this case). $K_{\max}$ values are included in Table 1.
E.5 Color-conditioning experiments
Table 4 gives $K_{\max}$ values and count distributions for the color-conditioning experiments of Figure 13. Unlike the location-conditioning experiments which use an automated classifier, the color experiment $K_{\max}$ values were evaluated by manually counting over 64 samples per configuration. We use the same $K_{\max}$ definition as in Section E.2 but with a 25% success threshold (generating $K$ objects at least 25% of the time, and at least K2K-2 objects at least 90% of the time), with objects appearing with approximately correct colors and with acceptable image quality. We test on up to $K=8$ colors.
Experiment $K_{\max}$ Manual counts
0 1 2 3 4 5 6 7 8 9 10
Color, trained 1 1 0 64 0 0 0 0 0 0 0 0 0
Color, trained 1-3 4 0 0 2 29 33 0 0 0 0 0 0
Color, trained 1-5 7 0 0 0 0 0 6 42 16 0 0 0
Table 4: $K_{\max}$ values and count distributions for color-conditioning experiments (Figure 13). Unlike the location-conditioning experiments which use an automated classifier, $K_{\max}$ is evaluated here by manual counting over 64 samples, since correct color attribution cannot be reliably assessed by an object-count classifier alone. Columns after the double line represent number of objects generated with correct colors (counts are evaluated at $K=K_{\max}$).
@@ Appendix F Pixel- and Conditional-Locality Gradient Analyses
In this section we describe the pixel- and conditional-locality analyses used in Figures 2, 2, 4, 11, 12.
Pixel-locality is measured as pixel gradient magnitude (average absolute of Jacobian from one pixel to all other pixels), and conditional locality is measured as conditional gradient magnitude (with the “gradient” estimated via a finite difference of the score computed with and without each conditioner). Further detail follows.
@@
$G_{i,j}(x,y,t)=\frac{1}{B}\sum_{b=1}^{B}\sum_{c=1}^{C}\left|\frac{\partial\widehat{x}_{0}[b,c,y,x]}{\partial x_{t}[b,c,i,j]}\right|$ (11)
@@
$I_{k}(x,y,t)=\left\|f_{\theta}(x_{t},t,C)[x,y]-f_{\theta}(x_{t},t,C_{-k})[x,y]\right\|_{1}$ (12)
@@
In Figure 12 we directly plot the gradient influence $I_{k}(x,y)$ for each conditioner $k$ across all times $t$. We quantify spatial locality by finding the smallest square centered at each target location that contains a specified fraction (90%) of the total gradient energy, i.e. the smallest square s.t. $\sum_{(i,j)\in\text{square}}G_{i,j}^{2}\geq 0.9\times\sum_{i,j}G_{i,j}^{2}$.
In Figure 2, Conditional Locality is an aggregated locality metric obtained by calculating the conditional gradient magnitude at each conditioned location, and computing
$\text{Conditional locality}=\frac{\sum_{k}I_{k}(x_{k},y_{k},t)}{\sum_{k}\sum_{k^{\prime}}I_{k}(x_{k^{\prime}},y_{k^{\prime}},t)}.$
@@
G.1 Pixel-space locality
For Figure 4 we use a pretrained SDXL model (out-of-the-box, no finetuning), specifically stabilityai/stable-diffusion-xl-base-1.0, with the prompt “a beautiful photograph with a horse in the middle, a dog on the left, and a cat on the right” – which contains implicit location conditioning. We perform a locality analysis similar to the one we used for CLEVR as described in Appendix F, but with a few adaptations. The gradient computation is performed in the VAE latent space. For conditional influence, we use a token-replacement ablation: for each non-stop-word in the prompt, we replace its token embedding(s) with the padding token embedding while keeping the sequence length fixed, then measure the change in the U-net prediction. This avoids re-encoding a modified prompt string and ensures the only difference is the ablated token. The difference is computed in float32 precision (even though the U-net runs in float16) to preserve small but meaningful spatial variations. All results are averaged over 10 seeds with deterministic noise for reproducibility. We ran the model with 50 inference steps using the standard EulerDiscreteScheduler. For each analysis timestep, we add noise to clean latents generated by the full pipeline (with classifier-free guidance) and perform a single-step U-net comparison, computing pixel Jacobians via autograd and word influence via the ablation described above. Results are shown at timesteps $t=1,25,50$ (low, mid, high noise). Additional single-seed results, alternative prompts, and spatial word-influence heatmaps are shown in Section G.1.1.
G.1.1 Additional pixel-space locality results
We present additional pixel-space locality results for SDXL across different seeds and prompts. Figure 14 extends the main-text analysis (Figure 4) to additional seeds and prompts. Figure 15 shows spatial word-influence heatmaps at low noise ($t=1$) and high noise ($t=50$) for two seeds of the three-object prompt used in the main text, as well as the swapped version (cat and dog positions exchanged). While some weak spatial localization of word influence is occasionally visible in individual heatmaps, this effect is subtle and inconsistent across seeds, prompts, and noise levels. Overall, the quantitative per-word influence profiles over time (Figure 4, right) are nearly identical at different pixel locations, confirming the absence of robust conditional-locality in pixel-space. This motivates the feature-space analysis in Section G.2.
Figure 14: SDXL pixel-space locality across seeds and prompts. (Top) All 10 seeds used for the main-text analysis (Figure 4) of the prompt “a beautiful photograph with a horse in the middle, a dog on the left, and a cat on the right.” (Bottom) Each row shows, for a given prompt and seed: the generated image (left column), pixel gradient magnitude heatmaps at low, mid, and high noise for a single pixel location (center columns, showing pixel-locality via localized cross patterns), and per-word conditional influence (ablation) over time at two pixel locations (right columns). Rows 1–3 are single-seed results for the three-object prompt (two seeds) and its swapped variant; rows 4–5 are 10-seed averages for attribute prompts (“a dog wearing a hat” and “a dog eating pizza”) which lack explicit location words. Across all seeds and prompts, the word-influence curves at different pixel locations are nearly identical, confirming the absence of conditional-locality in pixel-space.
Seed 42: “…horse in the middle, a dog on the left, and a cat on the right”
Seed 44: “…horse in the middle, a dog on the left, and a cat on the right”
Seed 42: “…horse in the middle, a cat on the left, and a dog on the right” (swap)
Figure 15: Spatial word-influence heatmaps in SDXL (single seeds). For each prompt and seed, word influence (token-replacement ablation) is shown at $t=1$ (low noise) and $t=50$ (high noise). Heatmaps use a shared colorscale across words within each timestep. Some weak spatial localization of word influence is occasionally visible, but it is not consistent across seeds, prompts, or noise levels. Seed 42, prompt 0 corresponds to a single seed of the 10-seed average shown in Figure 4.
G.2 Feature-space disentanglement